### v1

# Finiteness criteria for the solutions of the sequence of semi-decomposable form equations and inequalities

Abstract.
In this paper, we give some finiteness criteria for the solutions of the sequence of semi-decomposable form equations and inequalities, where the semi-decomposable form is factorized into a family of homogeneous polynomials with the distributive constant exceeding a certain number.
footnotetext: 2010 Mathematics Subject Classification: Primary 11J68; Secondary 11J25, 11J97.
   Key words and phrases: Diophantine approximation; subspace theorem; homogeneous polynomial.
1. Introduction
Let $k$ be a number field and let $M_{k}$ be the set of places. Let $S$ be a subset of $M_{k}$ which contains all archimedean places of $k$. Denote by $\mathcal{O}_{S}$ and $O_{S}^{*}$ the set of all $S$-integers and the set of all $S$-units, respectively. Denote by $H_{S}({\bf x})$ the $S$-height of the point ${\bf x}\in k^{m+1}$ (see Section 2 for detail definitions).
Let $F(x_{1},\ldots,x_{m})$ be a decomposable form of degree $q$ which can be factorized into linear factors over a finite extension $k^{\prime}$ of $k$. An interesting problem in the diophantine approximation is to consider the solutions of the following decomposable form inequations (see [5, 10, 11, 12] for reference):
$0<\prod_{v\in S}\|F(x_{0},\ldots,x_{m})\|_{v}\leqslant cH_{S}^{\lambda}(x_{0},\ldots,x_{m})\quad\text{ in }\quad(x_{0},\ldots,x_{m})\in\mathcal{O}_{S}^{m+1},$
where $S$ is a finite set of places of $k$ containing the archimedean places of $k$, $c$ and $\lambda$ are two given positive real numbers. If $(x_{0},\ldots,x_{m})$ is a solution of the above inequality, then so is $(\eta x_{0},\ldots,\eta x_{m})$ for every $\eta\in\mathcal{O}_{S}^{*}$. The solution $(\eta x_{0},\ldots,\eta x_{m})$ is said to be $\mathcal{O}_{S}^{*}$-proportional to $(x_{0},\ldots,x_{m})$.
In [3], Györy and Ru proved the following result for a sequence of decomposable form inequalities.
Theorem A (see [3, Theorem 2]). Let $q,m$ be positive integers. Let $c,\lambda$ be real numbers with $c>0,\lambda<q-2m$, and $k^{\prime}$ a finite extension of $k$. For $n=1,2,\ldots$, let $F_{n}({\bf x})=F_{n}(x_{0},\ldots,x_{m})\in\mathcal{O}_{S}[{\bf x}]$ denote a decomposable form of degree $q$ which can be factorized into linear factors over $k^{\prime}$, and suppose that these factors are in general position for each $n$, which means if we consider the hyperplanes in $\mathbb{P}^{m}$ defined by these linear factors, they are in general position in $\mathbb{P}^{m}$. Then there does not exist an infinite sequence of $\mathcal{O}_{S}^{*}$-nonproportional ${\bf x}_{n}\in\mathcal{O}_{S}^{m+1},n=1,2,\ldots$, for which
$0<\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}\leqslant cH_{S}^{\lambda}({\bf x}_{n})$
and $h(F_{n})=o(h({\bf x}_{n}))\text{ as }n\rightarrow\infty.$
In [9], by using the methods of Corvaja-Zannier [2] and Evertse-Ferretti [6, 7] and by introducing the notion of “distributive constant” of an arbitrary family of hypersurfaces with respect to a projective variety the author gave a general Schmidt’s subspace theorem. Motivated by these works, in this paper we will use the the notion of “distributive constant” to improve and generalize the above result to the case where each $F_{n}$ is a semi-decomposable form with a certain distributive constant. Here, a homogeneous polynomial $F(x_{0},\ldots,x_{m})\in k[x_{0},\ldots,x_{m}]$ is said to be semi-decomposable with the distributive constant $\Delta$ if $F$ can be factorized into a product of homogeneous polynomials $Q_{1},\ldots,Q_{q}$ over $\bar{k}$ with the distributive constant $\Delta$
In oder to state our result, we recall the following.
Definition 1.1.
Let $V$ be a projective subvariety of ${\mathbb{P}}^{m}$ defined over a number field $k$, and let $Q_{1},\ldots,Q_{q}$ be $q$ homogeneous polynomials in $\bar{k}[x_{0},\ldots,x_{m}]$. We define the distributive constant of the family $\mathcal{Q}=\{Q_{1},\ldots,Q_{q}\}$ with respect to $V(\bar{k})$ by
$\Delta_{\mathcal{Q},V(\bar{k})}:={\mathrel{\mathop{{\max}}\limits_{{\emptyset\neq\Gamma\subset\{1,\ldots,q\}}}}}\dfrac{\sharp\Gamma}{\dim V(\bar{k})-\dim(\bigcap_{j\in\Gamma}Q_{j}^{*})\cap V(\bar{k})}.$
Throughout this paper, if $V={\mathbb{P}}^{m}$ then we just write $\Delta_{\mathcal{Q}}$ for $\Delta_{\mathcal{Q},{\mathbb{P}}^{m}(\bar{k})}$ and call it the distributive constant of $\mathcal{Q}$.
Here we note that $\dim\emptyset=-\infty$, by $\sharp\Gamma$ we denote the number of elements of the set $\Gamma$ and by $Q^{*}$ we denote the support of the homogeneous polynomial $Q$, ie.,
$Q^{*}=\{(x_{0}:\cdots:x_{m})\in{\mathbb{P}}^{m}(\bar{k})|Q(x_{0},\ldots,x_{m})=0\}.$
Remark 1.2.
(a) If $V^{\prime}$ is a subvariety of $V$ and $Q_{j}\not\subset V^{\prime}$ for all $j\in\{1,\ldots,q\}$ then
$\Delta_{\mathcal{Q},V^{\prime}(\bar{k})}\leqslant(\dim V(\bar{k})-\dim V^{\prime}(\bar{k})+1)\Delta_{\mathcal{Q},V(\bar{k})}.$
Indeed, since $V^{\prime}\subset V$ and $Q_{j}\not\subset V^{\prime}$ for all $j\in\{1,\ldots,q\}$, then for every nonempty subset $\Gamma$ of $\{1,\ldots,q\}$,
$(\dim V(\bar{k})-\dim V^{\prime}(\bar{k})+1)(\dim V^{\prime}(\bar{k})-\dim V^{\prime}(\bar{k})\cap\bigcap_{j\in\Gamma}Q_{j}^{*})$
$\geq\dim V(\bar{k})-\dim V^{\prime}(\bar{k})+\dim V^{\prime}(\bar{k})-\dim V^{\prime}(\bar{k})\cap\bigcap_{j\in\Gamma}Q_{j}^{*}$
$\geq\dim V(\bar{k})-\dim V(\bar{k})\cap\bigcap_{j\in\Gamma}Q_{j}^{*},$
i.e.,
$\frac{(\dim V(\bar{k})-\dim V^{\prime}(\bar{k})+1)\cdot\sharp\Gamma}{\dim V(\bar{k})-\dim V(\bar{k})\cap\bigcap_{j\in\Gamma}Q_{j}^{*}}\geq\frac{\sharp\Gamma}{\dim V^{\prime}(\bar{k})-\dim V^{\prime}(\bar{k})\cap\bigcap_{j\in\Gamma}Q_{j}^{*}}.$
This straightforwardly implies that $\Delta_{\mathcal{Q},V^{\prime}(\bar{k})}\leqslant(\dim V(\bar{k})-\dim V^{\prime}(\bar{k})+1)\Delta_{\mathcal{Q},V(\bar{k})}.$
(b) By the definition we also have $\Delta_{\mathcal{Q},V(\bar{k})}\geq 1$. If $Q_{1},\ldots,Q_{q}\ (q\geqslant\ell+1)$ are in $\ell-$subgeneral position with respect to $V$ then $\Delta_{\mathcal{Q},V(\bar{k})}\leq\ell-m+1$.
Our main result is stated as follows.
Theorem 1.3.
Let $\ell,m,d$ be positive integers. Let $k^{\prime}$ be a finite extension of $k$ and $S\subset$ $M_{k}$ be a finite set containing all archimedean places. For $n=1,2,\ldots$, let $F_{n}({\bf x})=F_{n}(x_{0},\ldots,x_{m})\in\mathcal{O}_{S}[{\bf x}]$ be a sequence of semi-decomposable forms of degree $\ell$ with the distributive constant not exceeding a positive number $\Delta$ for each $n$, where $F_{n}=Q_{1,n}\cdots Q_{q,n}$ over $k^{\prime}$ with $\deg Q_{j,n}=d_{j,n}\ (1\leqslant j\leqslant q)$ such that $\max_{1\leqslant\jmath\leqslant q}d_{j,n}\leqslant d$ for each $n$ and $\Delta_{\{Q_{j,n}\}_{1\leq j\leq q}}\leq\Delta$ with $\ell>d\Delta\left(\frac{m}{2}+1\right)^{2}$. Let $c,\lambda$ be real numbers with $c>0,\lambda<\ell-d\Delta\left(\frac{m}{2}+1\right)^{2}$. Then, there does not exist an infinite sequence of $\mathcal{O}_{S}^{*}$-non-proportional $x_{n}\in\mathcal{O}_{S}^{m+1},n=1,2,\ldots$, for which
(1.4) $0<\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}\leqslant cH_{S}^{\lambda}({\bf x}_{n})$
and
(1.5) $h(F_{n})=o(h({\bf x}_{n}))\text{ as }n\rightarrow\infty$
From the above theorem, we immediately get the following corollary.
Corollary 1.6.
Let $k$ be a number field and $\{Q_{1},\ldots,Q_{q}\}$ a family of $q$ homogeneous polynomials in variables $(x_{1},\ldots,x_{m})$ with coefficients in $k$ and the distributive constant not exceeding a positive number $\Delta$. Let $d_{j}:=\deg Q_{j}=\ (1\leq j\leq q),d:=\max_{1\leqslant j\leqslant q}d_{j}$ and $F=Q_{1}\cdots Q_{q}$. Assume that $\deg F>d\Delta(\frac{m}{2}+1)^{2}$. Then, for every finite set $S$ of places of $k$ containing the archimedean places of $k$, for each positive number $\lambda<\deg F-d\Delta(\frac{m}{2}+1)^{2}$ and for each constant $c>0$, the inequality
$0<\prod_{v\in S}\|F(x_{1},\ldots,x_{m})\|_{v}\leqslant cH_{S}^{\lambda}(x_{1},\ldots,x_{m})\quad\text{ in }\quad(x_{1},\ldots,x_{m})\in\mathcal{O}_{S}^{m}$
has only finitely many $\mathcal{O}_{S}^{*}$-non-proportional solutions.
This corollary is an improvement and also is an generalization of Theorem 4.2 in [1].
2. Auxiliary result
Let $k$ be a number field. Denote by $M_{k}$ the set of places (equivalence classes of absolute values) of $k$ and by $M^{\infty}_{k}$ the set of archimedean places of $k$. For each $v\in M_{k}$, we choose the normalized absolute value $|\cdot|_{v}$ such that $|\cdot|_{v}=|\cdot|$ on $\mathbb{Q}$ (the standard absolute value) if $v$ is archimedean, and $|p|_{v}=p^{-1}$ if $v$ is non-archimedean and lies above the rational prime $p$. For each $v\in M_{k}$, denote by $k_{v}$ the completion of $k$ with respect to $v$ and set
$n_{v}:$ $=[k_{v}:\mathbb{Q}_{v}]/[k:\mathbb{Q}];$
[…]
Let $S$ be a finite subset of $M_{k}$, which contains $M^{\infty}_{k}$. An element $x\in k$ is said to be an $S$-integer if $\|x\|_{v}\leqslant 1$ for each $v\in M_{k}\setminus S$. We denote by $\mathcal{O}_{S}$ the set of all $S$-integers. Each unit of $\mathcal{O}_{S}$ is called an $S$-unit. The set of all $S$-units forms a multiplicative group which is denoted by $\mathcal{O}_{S}^{*}$.
Throughout this paper, the number $m$ is fixed. For ${\bf x}=(x_{0},\ldots,x_{m})\in k^{m+1}$, define
$\|{\bf x}\|_{v}:=\max\{\|x_{0}\|_{v},\ldots,\|x_{m}\|_{v}\},\ v\in M_{k}.$
Let ${\bf x}=(x_{0}:\cdots:x_{m})\in{\mathbb{P}}^{m}(k)$. The absolute logarithmic height of ${\bf x}$ is defined by
$h({\bf x}):=\sum_{v\in M_{k}}\log\|{\bf x}\|_{v}.$
If $x\in k^{*}$, we define the absolute logarithmic height of $x$ by
$h(x):=\sum_{v\in M_{k}}\log^{+}\|x\|_{v},$
where $\log^{+}a=\log\max\{1,a\}.$ For ${\bf x}=(x_{1},\ldots,x_{m})\in k^{m}$, we also define the $S$-height as
$H_{S}({\bf x})=\prod_{v\in S}\|x\|_{v}$
and the logarithmic $S$-height as $h_{S}({\bf x})=\log H_{S}({\bf x})$. If ${\bf x}\in\mathcal{O}_{S}^{m}\setminus\{0\}$, then $H_{S}({\bf x})\geq 1$ and $H_{S}(\alpha{\bf x})=H_{S}({\bf x})$ for all $\alpha\in\mathcal{O}_{S}^{*}$. It is clear that $H_{S}({\bf x})\geq h({\bf x})$ for any ${\bf x}\in\mathcal{O}_{S}^{m}\setminus\{0\}.$
Let =$\mathbb{N}=\{0,1,2,\ldots\}$ and for a positive integer $d$, we set
$\mathcal{T}_{d}:=\{(i_{0},\ldots,i_{m})\in\mathbb{N}^{m+1}\ :\ i_{0}+\cdots+i_{m}=d\}.$
Let $Q=\sum_{I\in\mathcal{T}_{d}}a_{I}{\bf x}^{I}$ be a homogeneous polynomial of degree $d$ in $k[x_{0},\ldots,x_{m}]$, where ${\bf x}^{I}=x^{i_{0}}_{0}\ldots x^{i_{m}}_{m}$ for ${\bf x}=(x_{0},\ldots,x_{m})$ and $I=(i_{0},\ldots,i_{m})$. Define $\|Q\|_{v}=\max\{\|a_{I}\|_{v};I\in\mathcal{T}_{d}\}$. For each ${\bf x}=(x_{0},\ldots,x_{m})$, it is easy to see that
$\begin{cases}\|Q({\bf x})\|_{v}\leqslant\sharp\mathcal{T}_{d}\cdot\|Q\|_{v}\cdot\|{\bf x}\|_{v}&\text{ if }v\in M^{\infty}_{k},\\ \|Q({\bf x})\|_{v}\leqslant\|Q\|_{v}\cdot\|{\bf x}\|_{v}&\text{ if }v\in M_{k}\setminus M^{\infty}_{k}.\end{cases}$
The height of $Q$ is defined by
$h(Q)=\sum_{v\in M_{k}}\log\|Q\|_{v}.$
For each $v\in M_{k}$, we define the Weil function $\lambda_{Q,v}$ by
$\lambda_{Q,v}({\bf x}):=\log\frac{\|{\bf x}\|_{v}^{d}\cdot\|Q\|_{v}}{\|Q({\bf x})\|_{v}},\ {\bf x}\in{\mathbb{P}}^{m}(k)\setminus\{Q=0\}.$
Theorem 2.1 (see [9, Theorem 1.3]).
Let $k$ be a number field, $S$ be a finite set of places of $k$ and let $V$ be an irreducible projective subvariety of ${\mathbb{P}}^{N}$ of dimension $m$ defined over $k$. Let $\{Q_{1},\ldots,Q_{q}\}$ be a family of $q$ homogeneous polynomials of $\bar{k}[x_{0},\ldots,x_{N}]$ with the distributive constant $\Delta$ with respect to $V$. Then for each $\epsilon>0$,
$\sum_{v\in S}\sum_{j=1}^{q}\dfrac{\lambda_{Q_{j},v}({\bf x})}{\deg Q_{j}}\leqslant(\Delta(m+1)+\epsilon)h({\bf x})$
for all ${\bf x}\in V(\bar{k})$ outside a union of closed proper subvarieties of $V(\bar{k})$.
3. Semi-decomposable form equations and inequalities
In order to prove Theorem 1.3, we first prove the following version of the Schmidt’s subspace theorem.
Theorem 3.1.
Let $k$ be a number field, $S$ a finite set of places of $k$ and $V$ an irreducible projective subvariety of ${\mathbb{P}}^{N}$ of dimension $m$ defined over $k$. Let $\{Q_{1},\ldots,Q_{q}\}$ be a family of $q$ homogeneous polynomials in $k[x_{0},\ldots,x_{N}]$ with the distributive constant $\Delta$ with respect to $V(\bar{k})$. Then for each $\epsilon>0$,
$\sum_{j=1}^{q}\sum_{v\in S}\dfrac{\lambda_{Q_{j},v}({\bf x})}{\deg Q_{j}}\leqslant\left(\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon\right)h({\bf x})$
for all ${\bf x}\in V(\bar{k})\setminus\bigcup_{i=1}^{q}\{Q_{i}=0\}$ but finitely many points.
We note that the inequality in Theorem 3.1 is weaker the one in Theorem 2.1. However, the conclusion in Theorem 3.1 is that such an inequality holds for all but finitely many points, while in Theorem 2.1 the conclusion is that the inequality holds only outside a proper Zariski-closed set.
Proof.
We see that $(\frac{m}{2}+1)^{2}\geqslant m+1$. Then, by applying Theorem 2.1 for the family $\{Q_{1},\ldots,Q_{q}\}$, we conclude that the set of ${\bf x}\in V(\bar{k})\setminus\bigcup_{i=1}^{q}\{Q_{i}=0\}$ with
(3.2) $\sum_{j=1}^{q}\sum_{v\in S}\log\left(\frac{\|{\bf x}\|_{v}^{\deg Q_{j}}\cdot\|Q_{j}\|_{v}}{\|Q_{j}({\bf x})\|_{v}}\right)^{1/{\deg Q_{j}}}>\left(\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon\right)h({\bf x})$
is contained in a finite union of proper subvariaties of $V(\bar{k})$. Note that these subvarieties are of dimension at most $m-1$.
Take $V^{\prime}$ be one of them with $\dim V^{\prime}\geqslant 1$. Let $\Delta^{\prime}$ be the distributive constant of the family $\{Q_{1},\ldots,Q_{q}\}$ with respect to $V^{\prime}$. Applying Theorem 2.1 to $V^{\prime}$ with the family $\{Q_{1},\ldots,Q_{q}\}$, we get that the set of ${\bf x}\in V^{\prime}\setminus\bigcup_{j=1}^{q}\{Q_{j}=0\}$ with
(3.3) $\sum_{j=1}^{q}\sum_{v\in S}\log\left(\frac{\|{\bf x}\|_{v}^{\deg Q_{j}}\cdot\|Q_{j}\|_{v}}{\|Q_{j}({\bf x})\|_{v}}\right)^{1/\deg Q_{j}}>(\Delta^{\prime}(\dim V^{\prime}+1)+\epsilon)h({\bf x})$
is contained in a finite union of proper subvarieties of $V^{\prime}$. Of course these subvarieties are of dimension at least $m-2$. Note that, from Remark 3.8 one has $\Delta^{\prime}\leqslant(m-\dim V^{\prime}+1)\Delta$ and hence
$\Delta^{\prime}(\dim V^{\prime}+1)\leqslant\Delta(m-\dim V^{\prime}+1)(\dim V^{\prime}+1)\leqslant\Delta\left(\frac{m}{2}+1\right)^{2}.$
Hence, if a point ${\bf x}\in V^{\prime}\setminus\bigcup_{j=1}^{q}\{Q_{j}=0\}$ satisfies (3.2) then it must satisfy (3.3). Therefore, the set of ${\bf x}\in V^{\prime}\setminus\bigcup_{j=1}^{q}\{Q_{j}=0\}$ satisfying (3.3) is contained in a finite union of proper subvariaties of $V(\bar{k})$ of dimension at least $m-2$. This argument holds for every such subvariety $V^{\prime}$. Hence, the set of ${\bf x}\in V(\bar{k})\setminus\bigcup_{j=1}^{q}\{Q_{j}=0\}$ satisfying (3.3) is contained in a finite union of proper subvariaties of $V(\bar{k})$ of dimension at least $m-2$.
By this way, after $m-1$ steps, we conclude that the set of ${\bf x}\in V(\bar{k})\setminus\bigcup_{i=1}^{q}\{Q_{i}=0\}$ satisfying (3.3) is contained in a finite union of proper subvariaties of $V(\bar{k})$ of dimension zero, i.e, this set is finite. The theorem is proved. ∎
Proof of Theorem 1.3.
Let $S^{\prime}$ be a subset of $M_{k^{\prime}}$ which consists of the extension of the places of $S$ to $k^{\prime}$. Then every $S$-integer in $k$ is also an $S^{\prime}$-integer in $k^{\prime}$. Moreover, we have $H_{S}({\bf x}_{n})=H_{S^{\prime}}({\bf x}_{n})$ and
$\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}=\prod_{w\in S^{\prime}}\|F_{n}({\bf x}_{n})\|_{w}\text{ for }{\bf x}_{n}\in\mathcal{O}_{S}^{m+1}.$
So the inequality (1.4) is preserved when we work on $k^{\prime}$. Therefore, without loss of generality we may assume that $k^{\prime}=k$.
Suppose contrarily that there is an infinite sequence ${\bf x}_{n}=(x_{0,n},\ldots,x_{m,n})\in\mathcal{O}_{S}^{m+1}$ which satisfies (1.4) and (1.5). If the heights $\{h({\bf x}_{n})\}_{n\geqslant 1}$ are bounded then then the set $\{{\bf x}_{n};n\geqslant 1\}$ is actually a finite union of sets of proportional vectors. Therefore, we may suppose that $h({\bf x}_{n})$ are not bounded.
Without loss of generality, we may assume that $h({\bf x}_{n})\rightarrow\infty$ as $n\rightarrow\infty$. Then, by the assumption, (1.5) also holds. Furthermore it follows that $H_{S}({\bf x}_{n})\rightarrow\infty$ as $n\rightarrow\infty$ (since ${\bf x}_{n}\in\mathcal{O}_{S}^{m+1}$). By $\max_{0\leqslant j\leqslant m}(Q_{j,n})\leqslant h(F_{n})+O(1)$, from the inequality (1.5) we deduce that
$\max_{0\leqslant j\leqslant m}h(Q_{j,n})=o(h({\bf x}_{n}))\text{ as }n\rightarrow\infty.$
Take $\epsilon$ a positive number small enough so that $\ell>d\Delta\left(\frac{m}{2}+1\right)^{2}+\lambda+\epsilon$. By Theorem 3.1, there is an infinite subsequence ${\bf x}_{n_{k}}\in\mathcal{O}_{S}^{m+1},k=1,2,\ldots$, of $\{{\bf x}_{n}\}$, which without loss of generality we may assume to be $\{{\bf x}_{n}\}$ itself, such that
$\sum_{v\in S}\sum_{j=1}^{q}\dfrac{1}{d_{j,n}}\log\frac{\|{\bf x}_{n}\|_{v}^{d_{j,n}}\cdot\|Q_{j,n}\|_{v}}{\|Q_{j,n}({\bf x}_{n})\|_{v}}\leqslant\left(\Delta\left(\frac{m}{2}+1\right)^{2}+\frac{\epsilon}{d}\right)h({\bf x}_{n}).$
By the assumption that $d\geqslant\max_{1\leqslant j\leqslant n}d_{j,n}$, the above inequality implies that
(3.4) $\sum_{v\in S}\sum_{j=1}^{q}\log\frac{\|{\bf x}_{n}\|_{v}^{d_{j,n}}\cdot\|Q_{j,n}\|_{v}}{\|Q_{j,n}({\bf x}_{n})\|_{v}}\leqslant\left(d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon\right)h({\bf x}_{n}).$
On the other hand, since ${\bf x}_{n}\in\mathcal{O}_{S}^{m+1}$, we have
(3.5) $h({\bf x}_{n})\leqslant\log H_{S}({\bf x}_{n}).$
With the note that $\ell=\sum_{j=0}^{m}d_{j,n}$ for each $n$, from (3.4) and (3.5) we have
$\prod_{v\in S}\frac{\|{\bf x}_{n}\|_{v}^{l}\cdot\prod_{j=1}^{q}\|Q_{j,n}\|_{v}}{\|F_{n}({\bf x}_{n})\|_{v}}\leq(H_{S}({\bf x}_{n}))^{d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon}$
whence
$\frac{H_{S}^{l}({\bf x}_{n})\cdot\prod_{v\in S}\prod_{j=1}^{q}\|Q_{j,n}\|_{v}}{\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}}\leq(H_{S}({\bf x}_{n}))^{d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon}.$
It is clear that $\prod_{v\in S}\prod_{j=1}^{q}\|Q_{j,n}\|_{v}\geq c^{\prime}\prod_{v\in S}\|F_{n}\|_{v}\geq c^{\prime}$ for some positive constant $c^{\prime}$. Therefore, from (1.4), for $n=1,2,\ldots$, we have
$H_{S}^{l}({\bf x}_{n})$ $\leqslant\frac{1}{c^{\prime}}\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}\cdot(H_{S}({\bf x}_{n}))^{d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon}$
$\leqslant\frac{c}{c^{\prime}}(H_{S}({\bf x}_{n}))^{d\Delta\left(\frac{m}{2}+1\right)^{2}+\epsilon+\lambda}.$
Since $H_{S}({\bf x}_{n})\rightarrow\infty$ as $n\rightarrow\infty$, by letting $n\rightarrow\infty$ we get $\ell>d\Delta(\frac{m}{2}+1)^{2}+\lambda+\epsilon$. This is a contradiction. Hence, this completes the proof of Theorem 1.3. ∎
Remark 3.6.
a) In [8], Q. Ji and Q. Yan and G. Yu consider the case where each semi-decomposable form $F_{n}$ can be factorized into moving hypersurfaces in general position. Therefore, once we can define the notion of “distributive constant” for families of moving hypersurface with respect to a subvariety of ${\mathbb{P}}^{m}$ defined over a number field $k$ then we may generalize the result of Ji, Yan and Yu.
b) With the same arguments in [8] (see also [3]), we will get a result on semi-decomposable form equation as follows.
Consider the $S$-integer solutions of a sequence of semi-decomposable form equations of the forms
(3.7) $F_{n}({\bf x})=G_{n}({\bf x}),$
where $G_{n}({\bf x})$ are nonzero polynomials. Let ${\bf x}_{n}$ be the $\mathcal{O}_{S}^{*}$-nonproportional solutions of (3.7). There is a positive constant $c$ such that
$0<\prod_{v\in S}\|F_{n}({\bf x}_{n})\|_{v}=\prod_{v\in S}\|G_{n}({\bf x}_{n})\|_{v}\leqslant c\prod_{v\in S}\|G_{n}\|_{v}(H_{S}({\bf x}_{n}))^{\deg G_{n}}.$
Applying Theorem 1.3 for $c$ and $l-d\Delta\left(\frac{m}{2}+1\right)^{2}-1<\lambda<l-d\Delta(\frac{m}{2}+1)^{2}$, we get the following theorem.
Theorem 3.8.
Let $\ell,m$ be positive integers. Let $k^{\prime}$ be a finite extension of $k$ and $S\subset$ $M_{k}$ a finite set containing all archimedean places. For $n=1,2,\ldots$, let $F_{n}({\bf x})=$ $F_{n}(x_{0},\ldots,x_{m})\in\mathcal{O}_{S}[{\bf x}]$ denote a sequence of semi-decomposable forms of degree $\ell$ with the distributive constant not exceeding a positive number $\Delta$ for each $n$. Assume that $F_{n}=Q_{1,n}\cdots Q_{q,n}$ over $k^{\prime}$ with $\deg Q_{j,n}=d_{j},1\leqslant j\leqslant q$ and $\Delta_{\{Q_{j,n}\}_{1\leq j\leq q}}\leq\Delta$ for each $n$. Let $d=\max_{1\leqslant j\leqslant q}d_{j}$. Assume that $\ell>d\Delta\left(\frac{m}{2}+1\right)^{2}$. Let $G_{n}({\bf x})\in\mathcal{O}_{S}[{\bf x}]$ such that $\deg G_{n}<\ell-d\Delta\left(\frac{m}{2}+1\right)^{2}$ for each $n$. Then there does not exist an infinite sequence of $\mathcal{O}_{S}^{*}$-non-proportional ${\bf x}_{n}\in\mathcal{O}_{S}^{m+1}$ for which
$F_{n}({\bf x}_{n})$ $=G_{n}({\bf x}_{n})\neq 0,n=1,2,\ldots,$
$\log\prod_{v\in S}\|G_{n}\|_{v}$ $=o(\log H_{S}({\bf x}_{n}))\text{ as }n\rightarrow\infty,$
and $\text{and }h(F_{n})$ $=o(h({\bf x}_{n}))\text{ as }n\rightarrow\infty.$
In particular, for each nonzero $S$-integer $b$, the equation
$F(x_{0},\ldots,x_{m})=b.$
(under the given assumption in Corollary 1.6 for $F$) has finitely many $\mathcal{O}_{S}^{*}$-nonproportional solutions, provided $\deg F>d\Delta\left(\frac{m}{2}+1\right)^{2}$ with $d=\max_{1\leqslant j\leqslant q}d_{j}$.
Disclosure statement
No potential conflict of interest was reported by the author(s).